Home Insert Deaw Design Loyourt Reerences Malings eview View Hep Desin Layout Question 3 Suppose Professor Mary spends $1,000 per year on tax supplies and Cube Smart storage space. She recently received two job offers from a XYZ marketing firm- one offer were for $60,000 per year, and the other was for $50,000. However, she turned both jobs down to continue her tax preparation business and storage firm. If Donald sells 10 paintings per year at a price of $900 each: a. What is her total revenue? R-10 paintings $900 each- $9000 total revenue b. What are her accounting profits? Pacs $9,000-$1,000 (expenses)- $8,000 c. What are her economic profits? Show your work! d. Which ty pe of profit (accounting or economic profit) reflects the best use of he time? Explain TOS
Answer #1

Answer 3

a)

Total Revenue is the amount a firm or individual receives by selling its goods. Here she sold 10 Paintings at the rate of 900 each.

Hence Total Revenue = 900*10 = $9000

b)

It is given in the question that she spends $1000 on tax supplies and cube smart storage space. Hence her total explicit cost = $1000

Accounting Profit = Total Revenue - Explicit cost = 9000 - 1000 = 8000

Hence, Accounting Profit = $8000

c)

In order to calculate Economic Profit we have to subtract Implicit cost from the Accounting profit. Here implicit cost is the highest income she is sacrificing in order to start her new business. Here Highest income she is sacrificing is the XYZ's firm 1 Salary which equals $50,000

Hence,

Economic Profit = 8000 - 50000 = - $42000

d)

Best use of her time is best represented by her economic Profit. Because Economic profit helps us to understand that starting this business is profitable or not. If economic Profit is negative (like in this case) then this means that she can earn more if she did not start this business But Accounting profit does not consider alternative uses of this time. It tells us that she is enjoying profit of $8000 if she starts this business.

Hence Economic profit helps us to determine the best use of her time and indicates that she should not start this business and accept XYZ offer of $50000 salary.